Fermi National Laboratory


2004 Accelerator Update Archive

December 27, 2004-January 3, 2005
- Operations established five stores during the last week of 2004 and the beginning of 2005. Unfortunately the last store, which had our highest initial luminosity since the shutdown ended, 80E30, was lost due to the power outages on the Monday midnight shift.
- The Tevatron quenched two times just before stores were established during this period.

December 17-December 20
- During this 48 hour period, Operations established two stores that combined with an existing store provided approximately 19 hours and 48 minutes of luminosity to the experiments.
- TeV loses two stores moments after they were established and one store just before it was to be established.
- The Antiproton Source losses 4mA due to an emittance problem
- Separator sparks still plague the Tevatron
- The Antiproton Source suffered from Lithium lens trips

December 15-December 17
- During this 48 hour period Operations established two stores that provided approximately 18 hours and 51 minutes of luminosity to the experiments
- The TeV suffered a quench due to a separator spark
- Store 3853 had a luminosity of 54.6E30, which is the highest luminosity since shutdown

December 13-December 15
- During this 48 hour period, Operations established two stores. They provided approximately 21 hours and 54 minutes of luminosity for the experiments.
- The TeV suffered a quench due to a separator spark.
- TeV experts put the separators through a long conditioning mode.
- The Recycler stashed antiprotons.

December 10-December 13
- During the past 72 hours Operations has established one store that combined with an existing store supplied the experiments with approximately 33 hours and 44 minutes of luminosity.
- Switchyard experts established beam to Meson on Friday.
- Recycler stashed antiprotons.
- Just before establishing a store, the Tevatron quenched due to a separator problem.

December 8-December 10
- During the last 48 hours, Operations established two stores that delivered approximately 20 hours and 6 minutes of luminosity to the experiments.
- TeV separator sparks cause loss of luminosity and then loss of a store.
- Mouse causes safety trip in a Main Injector power supply.

December 6-December 8
- Operations put the TeV into shot setup on Monday, but the Main Injector had a problem with their phase detector.
- The Switchyard 120 beam problem was found.
- The I- Source died late on Monday night. The H- Source was brought on line early Tuesday.
- The TeV had a problem with an abort kicker.
- Operations established store 3821 with an initial luminosity of 10.03E30 on Tuesday afternoon. Unfortunately, a separator spark caused a quench three hours later.

December 3-December 6
- NuMI established beam to their absorber on Friday at 6:50 PM.
- The Tevatron suffers ramp dumps and quenches all weekend.
- Pbar had emittance problems, but later fixed their damper system, which solved the trouble.
- The Tevatron Electron Lens was declared operational early Monday morning.

December 1-December 3
- Switchyard received permission to take beam on 12/1/04.
- Pbar stacking.
- MiniBooNE taking beam.
- NuMI will try to begin taking beam today, 12/3/04.

November 29-December 1
- The Tevatron and Recycler circulated beam on Monday.
- The Tevatron suffered a quench on Tuesday.
- The Tevatron ramped beam to 980 GeV on Wednesday morning.
- The Antiproton Source began slowly stacking.

November 24-November 29
- CHL lost a nitrogen compressor due to a power loss.
- Main Injector experts established a 120 GeV beam on Saturday.
- Tevatron experts established a partial turn of beam on Sunday, but then quenched at D2.
- The Antiproton Source took reverse protons during the Monday midnight shift.

November 19-November 22
- D0 experimenters turned on the solenoid on early Saturday morning.
- Cryo techs put a "band aide" on the TeV A4 helium leak and then started the A4 cooldown on Saturday at 3:30 PM.
- An expert began conditioning the MI-52 slow extraction septa.

November 12-November 19
The shutdown ends on Wednesday, December 1st.
- Tevatron: All scheduled work completed
- Linac: Up and running
- Booster: Up and running
- Antiproton Source: The plan for replacing the AP2 beam stop has been developed
- Main Injector/Recycler: All the components for the electron cooling have been installed
- MiniBooNE: All rigging has been completed
- NuMI: The vacuum pump down has been completed
- Cryogenic System: An air to helium leak was discovered on the TeV sector A49 spool piece

November 18
During the shutdown, the Accelerator Update will offer a series on the history and operation of the laboratory's accelerator complex. The tevatron is the tenth in a series.

November 10
During the shutdown, the Accelerator Update will offer a series on the history and operation of the laboratory's accelerator complex. The main injector is the ninth in a series.

October 29-November 5
- Tevatron: Shimming 412 Dipoles - 100% complete
- Linac: Up and Running
- Booster: Up and Running
- Antiproton Source: Technicians expect to begin closing up the Prevault and Pretarget areas the week beginning 11/8/04.
- MI/Recycler: MI-30 & 62 correction magnet replacement is approximately 95% complete.
- MiniBooNE: They are waiting on parts, but they expect to be ready by November 30th.
- NuMI: They have successfully tested their horn.

October 29
During the shutdown, the Accelerator Update will offer a series on the history and operation of the laboratory's accelerator complex. The main injector history is the eighth in a series.

October 22-29
- Tevatron: D17 separator installation - 99% complete.
- Linac: Up and running.
- Booster: Hit 4E12 at 70% efficiency.
- Antiproton Source: The Kicker beam pipe coating was completed.
- Main Injector/Recycler: The overall MI-31 Electron Cooling is 69% complete.
- MiniBooNE: The Horn will be tested on October 29th.
- NuMI: Their horn has been successfully tested.

October 15-22
- Tevatron: The D17 Separator installation is 90% complete. Its bake begins 10/22/04
- Linac: Up and running beam
- Booster: Vacuum problems slowing startup, but they should have beam by the afternoon on 10/22/04
- Antiproton Source: Analysis of the D/A lines is in progress
- Main Injector/Recycler: The overall shutdown progress is at 80%
- MiniBoone: The new horn has been installed
- NuMI: Their new horn has been installed and its testing may begin near the end of the month

October 22
During the shutdown, the Accelerator Update will offer a series on the history and operation of the laboratory's accelerator complex. The booster is the seventh in a series.

October 8-15
-Tevatron: The Crystal collimator has been installed.
-Linac: Linac began sending beam to its dump.
-Booster: Booster expects to be up and running beam by 10/20/04.
-Antiproton Source: The septum magnet has been pulse tested at full current.
-Main Injector/Recycler: The overall shutdown work is about 80% complete.
-MiniBooNE: The old horn has been moved to TSB.
-NuMI: The kicker is pulsing

October 15
During the shutdown, the Accelerator Update will offer a series on the history and operation of the laboratory's accelerator complex. The booster history is the sixth in a series.

October 1-8
-Linac: All work should be completed and the machine ready for beam by October 12 or 13.
-Booster: All work should be completed and the machine ready for beam by October 14 or 15.
-Antiproton Source: The ARF1-1 installation is about 95% complete.
-Main Injector: The LCW system is up and running.
-Tevatron: The E2, F2, & F3 conning tower work has been complete.
-MiniBoone: The Horn replacement is behind schedule by one week due to many problems.

October 6
During the shutdown, the Accelerator Update will offer a series on the history and operation of the laboratory's accelerator complex. The linac side-coupled cavities is the fifth in a series.

October 1
During the shutdown, the Accelerator Update will offer a series on the history and operation of the laboratory's accelerator complex. The linac is the fourth in a series.

September 17-24
-Tevatron: The D17 separator installation is 50% complete. They discovered a vacuum leak.
-Linac: The replacement of drift-tube tank #5 is 15% complete.
-Antiproton Source: The ARF1-1 installation is about 90% complete. It had a successful leak check and its final assembly is in progress.
-Main Injector: The MI-10 cooling tank repairs are about 55% complete.
-NuMI: The Horn replacement is about 52% complete.

September 22
During the shutdown, the Accelerator Update will offer a series on the history and operation of the laboratory's accelerator complex. The transport lines is the third in a series.

September 10-17
-Tevatron A0 Vacuum Upgrade - 60%
-Antiprotons Source Debuncher Motorized Quadrupole installation - ~30%
-Recycler/Main Injector Cooling Tank Repairs - 45%
-Main Injector Kicker Survey - 99%

September 15
During the shutdown, the Accelerator Update will offer a series on the history and operation of the laboratory's accelerator complex. The pre-accelerator is the second in a series.

September 4-10
- Beam Position Monitor upgrade in the Tevatron tunnel is 100% complete
- In the Linac, the High Energy RF maintenance, the mod VME fans and the drift tube tank #1 vaccuum repairs are all 100% complete
- The rebuild of the area for the Debuncher injection area beampipe and quadrupole replacement in the Antiproton Source has begun.
- Vacuum modification at MI-62 is 100% complete

September 10
During the shutdown, Accelerator Update will offer a series on the history and operation of the laboratory's accelerator complex. The Cockcroft-Walton pre-accelerator leads off the series.

August 30-September 3
The accelerator complex is in its 2nd week of a 13-week shutdown.
Tevatron:
- Only four sectors warmed up to room temperature
- 20% of shimming of dipoles complete
Linac:
- Drift Tube Tank #1 Vacuum Repairs
Booster:
- Vacuum Work - 48% complete
Antiproton Source:
- Quads alignment complete

August 20-August 23
- During this 72 hour period Operations established two stores that contributed approximately 24 hours of beam to the experiments
- TeV experts reported that the ground fault disappeared
- Operations established store 3744 on Saturday night and store 3745 on Sunday afternoon
- Nearly all the machines conducted studies
- The TeV turned off at 4 AM
- Fermilab began its 13-week shut down at 7 AM

August 18-August 20
- During this 48 hour period Operations established no stores. An existing store provided approximately 25 hours and 1 minute of luminosity to the experiments.
- Many machine studies occurred during this time period.
- The TeV lost its ramp due to a ground fault on Thursday afternoon, 8/19/04. The search for this ground fault still continues.
- The Fermi accelerator complex begins its 13-week shutdown on Monday, August 23, 2004 at 7 AM.

August 16-August 18
- During this 48 hour period Operations established two stores that contributed approx. 27 hours and 22 minutes of luminosity to the experiments.
- Booster had many GMPS problems
- The first attempt to establish a store on Monday after TeV turn on was aborted due to emittance and coalescing problems.
- Slipstacking is now the normal mode of operations for Main Injector and Pbar.
- The TeV quenched during the next shot setup

August 13-August 16
- During this 72 hour period Operations established no stores and provided approximately 13 hours and 45 minutes of luminosity to the experiments.
- Kautz Road substation goes down

August 11-August 13
- During this 48 hour period Operations established one store that provided approximately 41 hours of luminosity to the experiments.
- TeV bulk power supply developed a water leak.

August 9-August 11
- During this 48 hour period Operations established one store that provided approximately 29 hours and 15 minutes of luminosity to the experiments.
- Pbar stack was lost
- Main Injector power supply problems

August 6-August 9
- During this 72 hour period Operations established two stores that provided approximately 50 hours and 45 minutes of luminosity to the experiments.
- Booster experiences vacuum problems.
- TeV has to change out a kautzky valve.

August 4-August 6
- During this 48 hour period Operations established one store that provided approximately 20 hours and 3 minutes of luminosity to the experiments.
- Pbar has stacking trouble
- TeV store 3711 established by Mix Antiproton Mode
- CHL struggles with unstable D1 cryo system
- Feeder 32 trips off

August 2-August 4
- During this 48 hour period Operations established one store that provided approximately 24 hours and 47 minutes of luminosity to the experiments.
- Power supply trip causes partial loss of Recycler stash
- MiniBooNE tests show that ground-faulted Horn can't be run
- TeV suffers quench during dry squeeze
- Pbar stack loss due to over cooling
- TeV vacuum problems require access
- Debuncher kicker trouble halts stacking
- Lightning causes many trips
- Feeder 47 looses two phases

July 30-August 2
- During this 72 hour period Operations established one store that, combined with an existing store, provided approximately 15 hours and 13 minutes of luminosity to the experiments.
- The TeV suffers two store ending quenches
- The TeV quenched twice more during a dry squeeze
- B0 quadrupole (B0Q3) developed ground fault
- A Linac RF station (LRF4) has LCW leak
- LRF5 power amplifier replaced

July 28-July 30
- During this 48 hour period Operations established one store that, combined with an existing store provided the experiments with approximately 37 hours and 38 minutes of luminosity.
- MI suffers from corrupt MECAR data table
- Recycler stashes antiprotons
- TeV quenches during dry squeeze
- Booster down due to a 480V breaker trip in the West Gallery Yard

July 26-July 28
- During this 48 hour period Operations established two stores that combined with an existing store provided the experiments with approximately 37 hours and 13 minutes of luminosity.
- Recycler and Main Injector experts conduct many TLG studies
- Main Injector experts conducted many other studies
- The MiniBooNE Horn required a lower rep rate to keep from tripping off

July 23-July 26
- During this 72 hour period Operations established two stores that provided the experiments with approximately 48 hours and 6 minutes of luminosity.
- MiniBooNE suffers from many horn trips
- Store 3671 raised from an initial luminosity of 28.9E30 to 62.2E30 by adjusting separator voltage and collimators
- Booster RF station #1 tripped off many times

July 21-July 23
- During this 48 hour period Operations established one store that combined with an existing store provided approximately 39 hours and 57 minutes of luminosity to the experiments.
- I- Source generator motor stops
- A Linac quadrupole (QPS133) having problems
- Antiprotons lost from stack
- TeV suffers quench on Friday midnight shift

July 19-July 21
- During this 48 hour period Operations established one store that, combined with an existing store, provided approximately 43 hours and 6 minutes of luminosity to the experiments.
- Booster RF has many station trips
- Linac has RF and quadrupole problems
- TeV suffers quench during end-of-store study
- Store 3663 established by mixed mode antiprotons
- I- Source tripped off due to motor generator trouble

July 16-July 19
- A new initial luminosity record of 103.43 E30.
- Operations established 3 stores during the last seventy-two hours. They delivered approximately sixty-one hours and forty-five minutes of luminosity to the experiments.

July 14-July 16
- During this 48 hour period Operations established one store that combined with an existing store provided approximately 43 hours and 15 minutes of luminosity to the experiments.
- TEL problems didn't affect store
- Recycler stashed antiprotons

July 12-July 14
- During this 48 hour period Operations established one store that, added to an existing store, provided the experiments with approximately 31 hours and 32 minutes of luminosity.
- Meson has LCW troubles
- Switchyard had FSEPV problems
- E1 cryo wet engine bearings replaced
- Recycler stashes antiprotons

July 9-July 12
- During this 72 hour period Operations established 3 stores that provided approximately 47 hours and 9 minutes of luminosity to the experiments.
- Pbar water leaks repaired
- Pbar Lithium lens trips off; pulse magnet ground fault
- ComEd power glitch cause the loss of store 3641; The stack and stash survives
- Recycler stashes antiprotons successfully on Friday, but not on Monday

July 7-July 9
- Operations established one store during this 48 hour period. It provided approximately 4 hours and 9 minutes of luminosity to the experiments.
- TeV quench loses store 3629
- Power supply HF49 causes TeV quench
- AQ0 problems delay establishing store

July 5-July 7
- During this 48 hour period Operations established one store that, added to an existing store, provided approximately 32 hours and 47 minutes of luminosity to the experiments.
- Linac beam suffers from LRF5 water problems
- Booster suffers from BRF12 problems
- Store 3621 set NEW LUMINOSITY RECORD with 92.1E30
- TeV store 3621 lost due to a quench

July 2-July 5
- During this 72-hour period Operations established 3 stores that provided the experiments with approximately 57 hours and 44 minutes of luminosity.
- Recycler stashes antiprotons after every store
- The TeV suffers D0 quench
- The TeV Loses power to E0 service building to compressor trip
- A Linac RF station (LRF5) has water pump problem

June 30-July 2
- During this 48 hour period Operations established two stores that provided approximately 20 hours and 38 minutes of luminosity to the experiments.
- TeV sextupole power supply problems causes abort and holds off HEP

June 28-June 30
- There were no stores established during this 48 hour period. An existing store provided approximately 24 hours and 2 minutes of luminosity to the experiments.
- The TeV suffers a quench just after all protons and antiprotons (from Recycler and Pbar) were loaded
- TeV sextupole magnet (T:SF) ground faults

June 25-June 28
- During this 72 hour period of time operations established three stores that, added to an existing store, provided approximately 51 hours and 53 minutes of luminosity to the experiments.
- The vibrations from two earthquakes hit Fermilab causing the loss of one store
- A power glitch trips feeder 31 and hits Fixed Target areas and CHL
- Recycler stashes antiprotons

June 23-June 25
- During this 48 hour period Operations established two stores that combined with an existing store provided approximately 34 hours and 14 minutes of luminosity to the experiments.
- A TeV quench ends store 3588
- A ComEd power glitch trips off equipment in Linac and Booster
- Booster has problems with BRF12

June 21-June 23
- During this 48 hour period Operations established one store that, added to an existing store, provided the experiments with approximately 44 hours and 40 minutes of luminosity.
- Tevatron expert made changes to the collimator sequencer that reduced the problems with the colliding beam that has been troubling CDF and D0.
- Store 3688 set a NEW RECORD with an initial luminosity of 84.78E30.
- The Recycler lost its stash due to a Corrector magnet tripping off

June 18-June 21
- During this 72 hour period operations established two stores. Those stores and an existing store provided approximately 65 hours and 37 minutes of luminosity to the experiments and the accelerator experts.
- Recycler stashed antiprotons three times
- Booster and Linac suffer from many RF trips
- On the Monday midnight shift the TeV set a NEW LUMINOSITY RECORD with a luminosity of 82.3E30

June 16-June 18
- Operations established one store during this 48 hour period. This store combined with an existing store provided approximately 23 hours and 42 minutes of luminosity to the experiments.
- TeV changes separator helix voltage
- Recycler stashed antiprotons
- MI, Pbar, and TeV experts conduct many studies
- D0 requires access to repair detector

June 14-June 16
- During this 48 hour period Operations established two stores that combined with an existing store provided approximately 34 hours and 23 minutes of luminosity to the experiments.
- The accelerators set a NEW LUMINOSITY RECORD with a luminosity of 78.33E30
- An Accumulator RF station (ARF1) required tunnel access to repair
- A TeV quench aborted store 3574
- TeV experts help resolve a colliding beam proton problem

June 11-June 14
- During this 72 hour period Operations established one store that added to an existing store provided approximately 43 hours and 49 minutes of luminosity to the experiments.
- MI suffers a ground fault on Friday morning
- Lightning causes TeV abort early Saturday morning
- The Booster passed a goal on Sunday by extracting 6.03E12 protons in one pulse

June 9-June 11
- During this 48 hour period Operations established two stores that provided approximately 35 hours and 31 minutes of luminosity to the experiments.
- Accelerator experts establish the first Mixed-Antiproton store on June 9
- Booster suffers from a failed PFN kicker cable
- CDF data taking diminished due to colliding beam problems
- Pbar stack lost

June 7-June 9
- Operations established one store during this 48 hour period that combined with an existing store provided approximately 42 hours and 21 minutes of luminosity to the experiments.
- TeV sets NEW initial luminosity record of 76.8446E30
- TeV registers mysterious beam spike on Tuesday morning
- Off site power glitch causes many devices to trip off. Store remained intact.
- TeV, MI, Pbar, and Recycler prepare for Mixed Mode antiproton transfers

June 4-June 7
- During this 72 hour period Operations established two stores that provided approximately 44 hours and 52 minutes of luminosity to the experiments.
- TeV store aborted, cause still under investigation
- Recycler stashed antiprotons
- Accelerator experts conduct studies for mixed mode antiproton transfers later this week

June 2-June 4
- During this 48 hour period Operations established one store that combined with an existing store provided approximately 42 hours and 7 minutes of luminosity to the experiments.
- TeV conducts studies during D3 wet engine repairs
- Recycler stashes antiprotons
- Stacking interrupted due to a feeder 71 trip
- Pbar stack lost due to LCW leak in power supply
- Pbar sets stack longevity record of 69 days, 10 hours, and 8 minutes

May 31-June 2
- Operations established one store during this 48 hour period that combined with an existing store provided the experiments with approximately 43 hours and 45 minutes of luminosity.
- An AP1 Lambertson power supply needed daily attention.
- Pbar suffers from clock problem.
- Linac troubled by rectifier with water leak.

May 28-May 31
- During this weekend Operations established two stores that provided the experiments with approximately 14 hours and 49 minutes.
- The Tevatron suffered from a storm, an abort, a quench, a collimator, a sextupole power supply, and Lambertson problems that delayed the first shot until Sunday afternoon.

May 26-May 28
- Record store in Tevatron!
- Power glitch ends Record store prematurely.
- Tevatron recovery almost complete with Holiday weekend to follow.

May 24-May 26
- Operations established one store during this period of time that, added to an existing store, provided approximately 31 hours and 14 minutes of luminosity to the experiments.
- The TeV set a new record with an initial luminosity of 68.92E30
- The Recycler resumes stashing antiprotons

May 21-May 24
- During this weekend, Operations established two stores that provided the experiments with approximately 34 hours and 2 minutes of luminosity.
- Many problems by the TeV, MI, Pbar, and Booster machines, plus CDF and D0, led to a long delay before Operations established the first weekend store.

May 19-May 21
- During this period of time Operations established two stores that added with an existing store provided approximately 33 hours and 56 minutes of luminosity to the experiments.
- The TeV suffers two aborts and one quench
- TeV beam gap problem resolved
- Booster troubled by power supply problems

May 17-May 19
- Operations established two stores during this period of time that, added to an existing store, provided the experiments with approximately 16 hours and 32 minutes of luminosity.
- Colliding beam problems continue to bother the TeV
- Linac has stuck vacuum valve
- C3 cold compressor trips off and causes TeV quench

May 14-May 17
- During this 72 hour period Operations established three stores. Those stores, along with an existing store, provided approximately 44 hours and 58 minutes of luminosity to the experiments.
- Booster had two serious problems with ORBMP
- Two TeV stores suffer from a multi-house quench and an abort due to a vacuum problem.
- Pbar stacking halted due to a target shutter getting stuck and due to a power supply problem.

May 12-May 14
- During this period of time Operations established two stores that provided approximately 23 hours and 45 minutes of luminosity to the experiments.
- Tevatron ramp dump terminated one store early.

May 10-May 12
- During this period of time Operations established one store that provided approximately 38 hours of luminosity to the experiments.
- The Main Injector suffers from power supply problems

May 7-May 10
- Accelerator runs well over the weekend.
- MI power supply problems on Monday put a damper on the next Tevatron store.

May 5-May 7
- During this period of time Operations established one store that added to an existing store provided approximately 47 hours and 12 minutes of luminosity to the experiments.
- The Main Injector suffers from Lambertson trouble

May 3-May 5
- Operations established one store during this 48 hour period, which added to the existing store provided approximately 33 hours and 57 minutes of luminosity to the experiments.
- A clogged heat exchanger caused a Linac RF station to trip and hold off all beam
- A Main Injector power supply loop problem cut off stacking

April 30-May 3
- Operations established two stores that along with an existing store provided 62 hours and 55 minutes of luminosity to the experiments.
- The Meson MTest experiment takes fast spill
- The Antiproton Source had emittance, timing, and stacking problems

April 28-April 30
- During the last forty-eight hours operations established two stores that provided approximately 31 hours and 45 minutes of luminosity to the experiments.
- The Main Injector suffered many ramp trips
- The TeV shot setup halted due to a cogging problem
- The H- Source gets new hydrogen bottle

April 26-April 28
- Although Operations established one store during this time period, it wasn't useful beam. The existing store provided approximately 28 hours and 50 minutes of luminosity to the experiments.
- Core emittance problems plague the TeV, Pbar, and the Recycler
- Experts conduct many studies in the TeV, Main Injector, and Booster
- TeV sector A1 cold compressor fails early Wednesday morning

April 23-April 26
- Over the weekend Operations established two stores that when added to an existing store provided the experiments approximately 56 hours and 20 minutes of luminosity.
- Operation had trouble extracting antiprotons from Pbar for TeV store
- Meson lost some beam time due to sump alarm
- Operations forced to abort TeV protons due to separator problem
- Linac beam down on Monday morning due to a problem with the Klystron RF water skid

April 21-April 23
- During this time period Operations established one store that, along with an existing store, provided approximately 15 hours and 9 Minutes of luminosity to the experiments.
- The Antiproton Source replaced a pulsed magnet (PMAG).
- The TeV suffers a quench due to a crate problem.

April 19-April 21
- During this period of time Operations established two stores that provided approximately 27 hours and 30 minutes of luminosity to the experiments.
- The TeV suffers a quench on a proton abort.
- A power glitch tripped off some devices but doesn't bother the stack or store
- Tornado warnings drive operators to basement

April 16-April 19
- Operations established three stores during this period of time. Those stores provided approximately 25 hours and 34 minutes of luminosity to the experiments.
- TeV suffers two quenches due to antiproton kicker problems.
- A TeV separator problem requires Operations to terminate store.

April 14-April 16
- Operations established one store during this period of time. That store combined with an existing store provide the experiments with approximately 15 hours and 48 minutes of luminosity.
- A problem developed with the some Meson SWICs and SEMs.
- The TeV quenched early Thursday morning
- A TeV kicker prefire caused the quench. The kicker posed a problem all the way through to Friday morning.

April 12-April 14
- During this period of time Operations established one store that delivered approximately 12 hours and 25 minutes of luminosity to the experiments.
- Main Injector has problem accelerating beam for a study
- Tevatron experts conduct many hours of study

April 9-April 12
- During this period Operations established two stores. Unfortunately, one store was short lived and never provided useful luminosity. An existing store plus the one good store provided approximately 24 hours and 13 minutes of luminosity to the experiments.
- The TeV suffered two quenches.
- Recycler takes shots of antiprotons from Pbar.

April 7-April 9
- During the time period Operations established one store. That store along with an existing store provided the experiments with approximately 43 hours and 59 minutes of luminosity.
- Booster expert repair collimator.
- MI RF pond water algae causes strainer problems

April 5-April 7
- During this period Operations established two stores, which provided the experiments with approximately 36 hours and 3 minutes of luminosity.
- Experts performed NuMI power supply and magnet testing.
- Recycler stashed antiprotons twice during this period.

April 2-April 5
- During this period Operations established one store that provided the experiments with approximately 27 hours and 55 minutes of luminosity.
- The TeV quench on Friday midnight shift was due to the D1 bulk power supply
- An nearly established store quenched on Sunday due to a kicker prefire
- A feeder 35 power fault interferes with CHL and Meson systems
- A TeV A3 wet engine requires new flywheel

March 31-April 2
- During this period Operations established two stores. Those stores combined with an existing store provide the experiments with approximately 35 hours and 27 minutes of luminosity.
- The Tevatron suffered a six-house quench and then a multi-house quench
- Technicians continued to work on Switchyard vacuum problems

March 29-March 31
- During this period Operations established one store. That store with an existing store, provided the experiments with approximately 35 hours and 14 minutes of luminosity to the experiments.
- Quadrupole problems trouble the Tevatron
- Switchyard Septa vacuum work begins

March 26-March 29
- During the period Operations established two stores that provided approximately 27 hours and 29 minutes of luminosity to the experiments.
- On Friday afternoon a Pbar power supply tripped off and aborted the stack.
- The TeV suffered three quenches during studies this weekend.

March 24-March 26
- ComEd power glitch hits Linac, Booster, Pbar. Pbar held on to its stack.
- Alignment technicians finish aligning TeV components in CDF.
- Recycler stashes antiprotons

March 22-March 24
- On Tuesday technicians completed work on TeV sector A4 and cooldown began
- A magnet located in TeV sector F0 (HV100) required repair
- Operations successfully hipotted the TeV Wednesday morning

March 19-March 22
- The vacuum went bad at TeV sector E4; a valve failed.
- A phase problem developed between LRF1 and LRF2
- MI Experts discover wiring problem with bend magnet
- TeV shutdown continues

March 17-March 19
- Today is the fifth day of shutdown
- Operations established beam in the Linac on Thursday morning
- Operations established beam in Booster on Thursday evening
- TeV experts reported the A4 vacuum leak had returned. Two magnets need replacing

March 15-March 17
- The Fermi accelerator complex is in shutdown
- A partial list of shutdown work

March 12-March 15
- During this 72 hour period Operations established two stores that, along with an existing store, provided approximately 56 hours and 48 minutes of luminosity to the experiments.
- The TeV suffered a quench while aborting a study beam
- Recycler stashed antiprotons
- A MCenter power supply suffers a LCW leak
- TeV begins its ten-day shutdown. The rest of the machines begin a five-day shutdown.

March 10-March 12
- During this 48 hour period Operations established one store. That store along with an existing store provided the experiments with approximately 18 hours and 1 minute of luminosity.
- The TeV suffers two quenches
- The Recycler losses its stash
- MCenter takes beam
- The ten-day shutdown begins on Monday at midnight

March 8-March 10
- Fault on Feeder #35 causes problems.
- Operations established 1 store during the last forty-eight hours. They delivered approximately twenty-five hours of luminosity to the experiments.

March 5-March 8
- During the last 72 hours Operations established 2 stores. Those stores along with an already existing store provided approximately 56 hours and 30 minutes of luminosity to the experiments.
- Tevatron suffers quench at the beginning of study period on Friday.
- D-zero detector noise problems have disappeared.
- Two Linac RF stations and the TEL tripped off possibly due to a small power glitch.
- Recycler lost its stash of antiprotons Sunday evening.

March 3-March 5
- Operations established one store, which added to an already existing store, provided approximately 44 hours and 33 minutes of luminosity to the experiments.

March 1-March 3
- During the last forty-eight hours, Operations established one store. That store along with one that already existed provided approximately 23 hours and 40 minutes of luminosity to the experiments.
- The TeV quenched early Tuesday morning while ramping up for a store.

February 25-February 27
- During this period, Operations established one store. This store along with one that had already existed provided approximately 21 hours and 41 minutes of luminosity to the experiments.
- The Meson Test experiment at MT6 gets some beam every day.
- TeV suffers a six-house quench on Thursday.
- TeV vacuum problems delays cryo cooldown at A4.

February 23-February 25
- During this period Operations established one store. That store along with one that already existed provided approximately 44 hours and 47 minutes of luminosity to the experiments.
- Booster had power supply problems with GMPS and CEPNE1.
- Pbar had a power supply problem with A:IB.
- Operations sent beam to experiment located at MT6 in the Meson Detector Building.

February 20-February 23
- Operations established three stores during this period. Those stores along with one already established provided approximately 59 hours and 32 minutes of luminosity to the experiments.
- Operations established store 3245 with antiprotons from the Recycler.
- A failed Antiproton Source Septa connector halts stacking for four hours.

February 18-February 20
- During this period of time Operations established one store. That store along with an already existing store provided approximately 37 hours and 3 minutes of luminosity to the experiments.
- Store 3245 set a new luminosity record with 63.25E30.
- Booster suffers LCW tunnel leak.
- Recycler added to their stash so they could establish the next store.

February 16-February 18
- During this time period Operations established one store. That store, along with one that already existed, provided approximately 41 hours and 13 minutes of luminosity to the experiments.
- A 480V breaker in the East Booster Gallery tripped off.
- Experts from the Tevatron, Main Injector, Recycler, Antiproton Source, Booster, Linac, and Switchyard conducted machines studies.

February 13-February 16
-Fault on Feeder #37 causes some problems, overall a pretty good weekend.
- Operations established 2 stores during the last seventy-two hours. They delivered approximately fifty two hours and thirty minutes of luminosity to the experiments.

February 11-February 13
- Operations established one store during this period of time. That store plus one that already existed provided approximately 42 hours and 39 minutes of luminosity to the experiments.
- A NuMI power supply trip glitches a feeder that trips off Recycler, which loses its stash.
- Booster suffers from many LCW problems.
- Antiproton Source quadrupole trip causes loss of stack.

February 9-February 11
- Operations established one store during this period. That store along with an existing store, provided approximately 44 hours and 3 minutes of luminosity to the experiments.
- The I- Source suffers from sparking and trips
- Recycler takes transfers of antiprotons

February 6-February 9
- Operations established two stores during this time period. Those two stores along with an already existing store provided approximately 40 hours and 24 minutes of luminosity to the experiments.
- TEL failure causes A0 kicker room breaker to trip off
- Both Linac and Booster suffered from many RF problems over the weekend

February 4-February 6
- During this period Operations established one store. That store with an already existing store provided the experiments with thirty-six hours and thirteen minutes of luminosity.
- Store 3219 set a new record with an initial luminosity of 58.93E30.
- Two ComEd power glitches hit Fermilab on Wednesday. No serious harm done.

February 2-February 4
- During this period of time Operations established two stores. These stores provided the experiments with approximately thirty-seven hours and 32 minutes of luminosity.
- The Fermi accelerators set two new luminosity records with both stores. The record is now 55.5E30.
- Switchyard experts successful in sending beam to the Switchyard dump.

January 30-February 2
- During this period Operations established two stores. Those stores provided approximately forty-six hours and 24 minutes of luminosity to the experiments. The best luminosity of the two stores was 51.20E30.
- A harmonic filter on the Master Substation failed to close on Friday.
- D0 repaired a water leak, but then discovered a power supply problem.

January 28-January 30
- During this period Operations established one store. That store along with an existing store provided approximately twenty-five hours and 13 minutes of luminosity to the experiments.
- 100 mA of antiprotons lost from stack on Thursday.
- The D0 experimenters discover a water leak that required them to open their detector to make repairs.

January 26-January 28
- During this period Operations established three stores. Those stores, along with one that already existed, provided approximately twenty-one hours and twenty-five minutes of luminosity to the experiments.
- A LCW pressure problem causes store to abort.
- TeV wet engine required overhaul.
- To save from quenching operators abort store due to Tevatron cryo vacuum problem.

January 23-January 26
- During this period Operations established three stores. Those three stores provided approximately forty-eight hours and four minutes of luminosity for the experiments.
- Recycler completes its first transfer of antiprotons for a Tevatron store.
- The Antiproton Source technicians repair failed Debuncher Septum connection.
- Booster experts repair a capacitor and SCR Bridge for the ORBMP power supply.

January 21-January 23
- During this time period Operations established one store. That store along with an existing store provided approximately forty-one hours and seven minutes of luminosity to the experiments.
- Meson feeders 30 and 31 have both faulted.
- Linac switches to the H- Source.
- The Tevatron suffered a quench early Friday morning. There was no beam in the machine.

January 19-January 21
- During this time period Operations established one store. That store combined with one that already existed provided twenty-nine hours of luminosity to the experiments.
- The Tevatron suffered a sector F4 quench on Tuesday.
- Feeder 30 tripped off and caused a problem with CHL and Meson.

January 16-January 19
- During this period of time Operations established three stores. Those stores along with one already established provided approximately 53 hours and 31 minutes of luminosity to the experiments.
- An Antiproton Source Debuncher kicker caused several problems on Friday and Saturday.
- Experts established beam to the Switchyard dump on Sunday.
- Store 3177 provided a total integrated luminosity of 2020 inverse nano-barns.

January 14-January 16
- Another Tev quench.
- Operations established 1 store during the last forty-eight hours. They delivered approximately seventeen hours and forty-five minutes of luminosity to the experiments.

January 12-January 14
- Tev quenches continue.
- Operations established 1 store during the last forty-eight hours. They delivered approximately eleven hours and fifteen minutes of luminosity to the experiments.

January 9-January 12
- Quenches in the Tev and problems with the lithium lens in Pbar make for slow progress this weekend.
- Operations established 1 store during the last seventy-two hours. They delivered approximately twenty hours and thirty minutes of luminosity to the experiments.

January 7-January 9
- No major problems, good steady running
- Shot #3132 set a new one store integrated luminosity of 2.546 pb-1.
- Operations established 1 store during the last forty-eight hours. They delivered approximately forty two hours and thirty minutes of luminosity to the experiments.

January 5-January 7
- No major problems, good steady running.
- Operations established 1 store during the last forty-eight hours. They delivered approximately forty three hours and thirty minutes of luminosity to the experiments.

December 29-January 5
- During this period of time Operations established six stores, which delivered luminosity for approximately 95 hours and 40 minutes.
- Booster accessed twice, first to repair a small LCW leak, and second to repair a RF cavity.
- Pbar accessed to repair a Septa.

[an error occurred while processing this directive]

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.


Return to Current Accelerator Update



last modified 1/3/2005   email Fermilab